Today (T=0), an investor purchased a nine year bond with an 8.0% coupon for $9,680. The bond has a face value of $10,000. In six months (T=0.5) interest rates have increased by 1.0% and the investor decides to sell the bond immediately after receiving the first coupon payment. What is the investor’s total gain (loss) on the bond?

HINT: Total Gain (Loss) = Price Change in Bond + Coupon

A. ($583) B. ($183) C. ($150) D. $190 E. $990
Answer #1

ytm=RATE(9*2,8%*10000/2,-9680,10000)*2=8.516%
Price at T=0.5: =PV((8.516%+1%)/2,8.5*2,8%*10000/2,10000)=9129.77

Total gain=9129.77+10000*8%/2-9680=-150.23

Hence, Option C